OpenStudy (anonymous):

The sides of a quadrilateral measure 12, 18, 20, and 16.The longest side of a similar quadrilateral measures 5. Find the measures of the remaining sides of this quadrilateral.

Directrix (directrix):

@Malikdbuseck Because the quadrilaterals are similar, the ratio of the longest side of the first quadrilateral to the longest side of the second quadrilateral is the scale factor. 20/5 = 4/1 or 4:1 is the scale factor. You have three equations to solve: 1) 20/5 = 18/z Cross multiply and solve for z. 2) 20/5 = 16/y 3) 20/5 = 12/x